Assignment 1: Memo to Management

You are working in the finance department of Space Sky Flight Ltd (SSF). The Company has spent $6.5 million in research and development over the past 12 months developing a drone capable to fix satellites to compete in the space industry. SSF’s directors now need to choose between three options for bringing this product to the market. These options are:

Option A: Manufacturing the product “in-house” and selling directly to the market

Your task

Your boss, SSF’s CFO Savanah Harley, has asked you to evaluate the three different options and draft a memo to the Board of Directors providing recommendations on the alternatives, along with supporting analysis.

Savanah has outlined the following three areas you need to cover in your memo:

  1. Analyse base case figures for the three options and using NPV as the investment decision rule;
  2. Provide recommendations based on the base-case analyses;
  3. Provide recommendations on further analyses and factors that should be considered prior to making a final decision on the three options (Note. You do NOT have to undertake any further financial analyses).

Further details for the various options are as follows:

Option A

Two months ago, SSF paid an external consultant $950,000 for a production plan and demand analysis. The consultant recommended producing and selling the product for five years only as technological innovation will likely render the market too competitive to be profitable enough after that time. Sales of the product are estimated as follows:

Year

Estimated sales

volume (000’s of units)

1

4

2

3.5

3

5.5

4

3

5

1.5

In the first year, it is estimated that the product will be sold for $110,000 per unit. However, the price will drop in the following three years to $80,000 per unit and fall again to $60,000 per unit in the final year of the project, reflecting the effects of anticipated competition and improving technology in the market.

Variable production costs are estimated to be $45,000 per unit for the entire life of the project.

Fixed production costs (excluding depreciation) are predicted to be $10.5 million per year and marketing costs will be $8 million per year.

Production will take place in factory space the company owns and currently rents to another business for

$7 million per year. Equipment costing $300 million will have to be purchased. This equipment will be depreciated for tax purposes using the prime cost method at a rate of 20% per annum. At the end of the project, the company expects to be able to sell the equipment for $75 million.

Investment in net working capital will also be required. It is estimated that accounts receivable will be 25% of sales, while inventory and accounts payable will each be 20% of variable and fixed production costs (excluding depreciation). This investment is required from the beginning of the project because credit sales, inventory stocks and purchases on trade credit will begin building up immediately. All accounts receivable will be collected, suppliers paid and inventories sold by the end of the project, thus the investment in net working capital will be returned at that point.
Answer #1
Assume, Tax Rate =30%
Required Return =18%(Considering it is a risky Project)
Amount spent on Research and Development and amount paid to external consultants are sunk costs
These costs are not relevant for this analysis
Present Value(PV) of Cash Flow:
(Cash Flow)/((1+i)^N)
i=discount rate=Required Return=18%=0.18
N=Year   of Cash Flow
Annual Depreciation =300million*20% $60,000,000
Before tax Salvage Value $75,000,000
After Tax Salvage Cash Flow =75 million*(1-0.3) $52,500,000
N Year 0 1 2 3 4 5
A Initial investment -$300,000,000
b Sales in Units                     4,000                     3,500                     5,500                       3,000                    1,500
c Sales Price per unit $110,000 $80,000 $80,000 $80,000 $60,000
d=b*c Annual Sales Revenue $440,000,000 $280,000,000 $440,000,000 $240,000,000 $90,000,000
e Variable Production Cost per unit $45,000 $45,000 $45,000 $45,000 $45,000
f=b*e Total Variable Production Costs $180,000,000 $157,500,000 $247,500,000 $135,000,000 $67,500,000
g=d-f Contribution Margin $260,000,000 $122,500,000 $192,500,000 $105,000,000 $22,500,000
h Fixed Production Costs $10,500,000 $10,500,000 $10,500,000 $10,500,000 $10,500,000
m Marketing Costs $8,000,000 $8,000,000 $8,000,000 $8,000,000 $8,000,000
i Depreciation expenses $60,000,000 $60,000,000 $60,000,000 $60,000,000 $60,000,000
j=g-h-i-m Before tax Operating Income $181,500,000 $44,000,000 $114,000,000 $26,500,000 -$56,000,000
k=j*(1-0.3) After Tax Operating Income $127,050,000 $30,800,000 $79,800,000 $18,550,000 -$39,200,000
l Add:Depreciation (Non Cash expense) $60,000,000 $60,000,000 $60,000,000 $60,000,000 $60,000,000
X=k+l Annual Operating Cash Flow $187,050,000 $90,800,000 $139,800,000 $78,550,000 $20,800,000
n Opportunity cost of factory space -$7,000,000 -$7,000,000 -$7,000,000 -$7,000,000 -$7,000,000
Working Capital:
.(1) Accounts Receivable (25% of Sales) $110,000,000 $70,000,000 $110,000,000 $60,000,000 $22,500,000 $0
.(2) Accounts Payable(20% of Variable and fixed Production Costs) $38,100,000 $33,600,000 $51,600,000 $29,100,000 $15,600,000                           -  
.(3)=.(1)-.(2) Working Capital Needed $71,900,000 $36,400,000 $58,400,000 $30,900,000 $6,900,000 $0
Y Working Capital Cash Flow -$71,900,000 $35,500,000 -$22,000,000 $27,500,000 $24,000,000 $6,900,000
Z Terminal Salvage Cash Flow $52,500,000
CF=A+X+n+Y+X Net Cash Flow -$371,900,000 $215,550,000 $61,800,000 $160,300,000 $95,550,000 $73,200,000
PV=CF/(1.18^N) Present Valure -$371,900,000 $182,669,492 $44,383,798 $97,563,529 $49,283,627 $31,996,395 $33,996,840
NPV=Sum of PV Net Present Value $33,996,840
Yes, we should undertake this project
NPV is positive
Final decision to be taken after Sensitivity analysis and Scenario Analysis